OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). Explain how please?

OpenStudy (anonymous):

y=mx+n is a line eq. m is slope if the lines are parallel then their slope is the same. y=7x+2 its slpoe is m=7 its parallel lines has the same slope which is 7 if we know one point and slpoe we can find the eq. m=(y-y1)/(x-x1) 7=(y-(-3))/(x-1) 7x-7=y+3 y=7x-10 you see second line has the same slope m=7